OpenStudy (anonymous):

Find a vector w with the same direction of u but twice as long as v. u = {0 4 2} v = {2 1 0}. I did some caculate, but I think I am wrong. I done the length of v is √5 , unit vector with the same direction of u is 0,2/√5,1/√5 , but I think if w has same direction and twice long as v (2*√5) it will be {0 4 2}, same as u, am I wrong??

OpenStudy (anonymous):

how about getting a unit vector in the same direction as u?

OpenStudy (anonymous):

- find magnitude of v and then double it - find unit vector of u - multiply the above two

OpenStudy (anonymous):

- magnitude of v and then double it \[magnitude ||v|| = \sqrt{2^2 + 1^2 + 0^2} = \sqrt{5}\] double it become \[2\sqrt{5}\] - find unit vector of u \[magnitude ||u|| = \sqrt{0^2 + 4^2 + 2^2} = 2\sqrt{5}\] so \[unit.vector = \left\{ 0, \frac{ 2 }{ \sqrt{5} }, \frac{ 1 }{ \sqrt{5} } \right\}\] - multiply the above two result is {0, 4, 2} same as u, am I wrong?

OpenStudy (anonymous):

when u multiply u get ( 0, 4/5, 2/5)

OpenStudy (anonymous):

you are correct in your solution, @ccclp